I need help with -3(7-3n)+2n=5(2n-4).
Answer by stanbon(75887)

-3(7-3n)+2n=5(2n-4).
---
-21+9n+2n = 10n-20
11n-21 = 10n-20
n = 1
-------------
Cheers,
Stan H.